OpenStudy (anonymous):

The value of Mr. Cox’s car x years after its purchase is given by the function, V(x) = 20,000(0.85)x. Approximately, what was the value of Mr. Cox’s car 5 years after its purchase?

OpenStudy (allank):

The thing you need to do is solve for V(5).

OpenStudy (anonymous):

so how do i do that @allank

OpenStudy (allank):

Replace the x with 5 in 20,000(0.85)x

OpenStudy (anonymous):

ok

OpenStudy (allank):

Cool. A quick example: f(x) = 2x+10 therefore f(5) = 2(5)+10 = 20

OpenStudy (anonymous):

ook